Find the measure of an angle if the measure of is complement is 53 degrees. Thank you for all the help and sorry if this has disturbed you in any way.
Found 2 solutions by drj, checkley77:
Answer by drj(1380)

Find the measure of an angle if the measure of is complement is 53 degrees.

Step 1. Complementary angles means two angles add up to 90 degrees.

Step 2. Let x be the angle.

Step 3. Let 90-x be its complement.

Step 4. Given 90-x=53 or x=90-53=37.

Step 5. ANSWER: The angle is 37 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J

Answer by checkley77(12844)

If these are two complemental angles then:
x+43=90
x=90-43
x=47 is the other complemental angle.